OpenStudy (anonymous):

What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4?

OpenStudy (anonymous):

hey its august wh are you doin homework

OpenStudy (anonymous):

i take summer classes so i can get ahead.

OpenStudy (amistre64):

formula for the sum of geo is:\[a_1\frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

a=11, R=-4, and n=6

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so a1=1-(-4)^6/1-(-4)?

OpenStudy (amistre64):

hmm, close a1(1-(-4)^6)/1-(-4) 11(1-(-4)^6)/(1+4)

OpenStudy (amistre64):

\[11\frac{1-(-4)^6}{1+4}\]

OpenStudy (anonymous):

okay, and this makes the sun?

OpenStudy (amistre64):

..... it seems right to me :)

OpenStudy (amistre64):

-9009; and since the terms are few in number we can double check with brute math if need be

OpenStudy (anonymous):

okay thanks!!!!!!

OpenStudy (amistre64):

11 + 11(-4) + 11(-4)^2+ 11(-4)^3+ 11(-4)^4+ 11(-4)^5 11( 1 - 4 + 16 - 64 + 256 - 1024) 11( 273 - 1092) 11( -819) -819 -819 ------- -9009